### v3

# Dynamical scaling study for the estimation of dynamical exponent $z$ of three-dimensional XY spin glass model

Abstract
To analyze the ±$\pm J$ XY spin-glass in three dimensions, we verified a method aimed at obtaining a high-precision dynamical exponent $z$ from the correlation length in the nonequilibrium relaxation process. The obtained $z$ yielded consistent and highly accurate results in previous studies for relatively well-studied models—specifically, the three-dimensional (3D) ferromagnetic Ising model and the 3D ±$\pm J$ Ising model. Building on these previous studies, we used this method and the dynamical scaling method to analyze the 3D ±$\pm J$ XY model and obtained highly precise critical temperatures and exponents. These findings support the spin chirality decoupling picture, explaining the experimental spin-glass phase transition.
I Introduction
Spin-glass (SG) systems have been the subject of extensive study for many years, serving as a prototype for random systems and finding wide-ranging applications in areas such as quantum annealing and information sciences. Despite this extensive research, numerous aspects of the nature of SG phase transitions remain poorly understood. One such aspect pertains to the applicability of a spin chirality decoupling picture in explaining experimental Heisenberg-like SG magnets[1] (i.e., whether $T_{\mathrm{CG}}>T_{\mathrm{SG}}>0$ or $T_{\mathrm{CG}}=T_{\mathrm{SG}}>0$).
The lack of progress in SG research is attributable to the substantial computational time required to achieve equilibrium states as a result of the slow dynamics arising from frustration in SG systems. One effective approach to overcoming this issue is the nonequilibrium relaxation (NER) method. A recent study using this method confirmed the SG universality in the ±$\pm J$ Ising model in three dimensions[2]. In the NER method, obtaining the SG universality expressed by critical exponents $\nu$ necessitates the precise determination of the dynamical critical exponent $z$. The authors of a previous study used the asymptotic behavior of the dynamical Binder parameter $g_{\mathrm{SG}}$ (expressed as $g_{\mathrm{SG}}\sim t^{d/z}$, where $t$ denotes time and $d$ represents the dimensionality of the system) at the critical temperature[3]. By extrapolating the time $t$ to infinity, the authors were able to evaluate $z$ with high precision. However, this method is not applicable when the dynamical Binder parameter $g_{\mathrm{SG}}$ does not exhibit algebraic divergence[4]. One such instance is observed in the Binder parameter for a chiral glass in the 3D ±$\pm J$ XY model. To confirm universality in such SG systems, it remains imperative to compute $z$ with high precision even in such cases.
Typically, $z$ is defined using the dynamical correlation length $\xi(t)$ at the critical temperature, defined as:
$\xi(t)\sim t^{1/z}.$ (1)
Thus, if $\xi(t)$ can be evaluated, then $z$ can be estimated. The second moment method is commonly used to obtain the “static” correlation length, and this approach has yielded significant results in equilibrium Monte Carlo simulations. However, Nakamura have pointed out that even applying the second moment method to estimate the “dynamical” correlation length $\xi(t)$ may not lead to a precise determination of $z$: Even in simple 3D ferromagnetic (FM) Ising models, the results often deviate substantially from reliable benchmark values[5]. This discrepancy might be attributable to the nontrivial question of whether the second moment correlation length can be applied in nonequilibrium relaxation processes[5].
Recently, a method has been devised to estimate $\xi(t)$ using scaling laws for the dynamical correlation function $f(r,t)$ in conjunction with Gaussian process regression[6, 7]. Using this approach to obtain $\xi(t)$, in conjunction with Eq. (1), Nakamura confirmed that $z$ for the Heisenberg SG model can be computed[7]. Hence, this method is expected to enable $z$ to be calculated for various models, including those encompassing SG systems. However, a detailed validation of this assertion remains largely unexplored.
This paper has two main objectives. The first objective is to investigate the critical phenomena in the ±$\pm J$ XY model, which is an SG system in which the Binder parameter exhibits no algebraic divergence[4]. This approach enables an analysis of critical phenomena in various SG systems, such as Heisenberg systems and gauge glass models. The second objective is to establish the reliability of the estimated $z$ through our study because a precise value of $z$ is needed to study the SG systems through dynamical scaling analysis. To establish the reliability of the estimated $z$, we apply this method to relatively well-studied models—the 3D FM Ising model and the 3D ±$\pm J$ Ising model—and compute $z$.
This paper is structured as follows. In Sec. 2, we describe the models, observables, and methods used to estimate the critical temperature and dynamical exponent. In Sec. 3, we validate the estimation method of the dynamical exponent by applying it to the 3D FM Ising model and the 3D ±J Ising model. In Sec. 4, we apply this method to the 3D ±J XY model and analyze its critical behavior using dynamical scaling. Finally, Sec. 5 provides a summary and discussion.
II Methods
II.1 Models and Observables for Estimating the Critical Temperature
In the present study, we mainly aimed to analyze the ±$\pm J$ XY model in three dimensions. The Hamiltonian can be expressed as
$H=-\sum_{\langle i,j\rangle}J_{ij}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}=-\sum_{\langle i,j\rangle}J_{ij}\cos(\theta_{i}-\theta_{j}).$ (2)
The interaction $J_{ij}$ takes one of two values, $+J$ or $-J$, with the same probability. The summation $\sum_{\langle i,j\rangle}$ is taken over all nearest-neighboring sites. The temperature $T$ is scaled by $J/k_{\mathrm{B}}$. Each spin $\boldsymbol{S}_{i}$ is characterized by an angle $\theta_{i}$ in the $XY$ plane. The linear lattice size is denoted by $L$. The total number of spins is $N=L^{2}(L-1)$, and a skewed periodic boundary condition is imposed.
We calculated the SG susceptibility $\chi_{\rm SG}$ for SG systems. It is defined as
$\chi_{\rm SG}=\frac{1}{N}\Bigg{[}\sum_{i,j}\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}\rangle^{2}\Bigg{]}_{J},$ (3)
[…]
$\langle\boldsymbol{S}_{i}\boldsymbol{S}_{j}\rangle=\frac{1}{n}\sum_{A=1}^{n}\boldsymbol{S}_{i}^{(A)}\boldsymbol{S}_{j}^{(A)},$ (4)
where $n$ denotes the replica number and $A$ denotes the replica index. The magnitude of $n$ affects the accuracy of thermal averages. Each replica system has the same bonds $J_{ij}$ but starts with a different randomly assigned initial spin state. The replicas’ overlap between $A$ and $B$ is defined by
$q_{\mu\nu}^{AB}=\frac{1}{N}\sum_{i}^{N}S_{i\mu}^{(A)}S_{j\nu}^{(B)}.$ (5)
where subscripts $\mu$ and $\nu$ denote the components of spin. The SG susceptibility can be rewritten using this overlap as
$\chi_{\mathrm{SG}}=\frac{N}{C_{n}}\Bigg{[}\sum_{A>B}\sum_{\mu,\nu}(q_{\mu\nu}^{AB})^{2}\Bigg{]}_{J}.$ (6)
Here, $C_{n}=n(n-1)/2$ represents the number of combinations for choosing two replicas out of $n$ ones.
The existence of the chiral-glass (CG) transition in the 3D ±$\pm J$ XY model has been confirmed in numerous previous studies[8, 4, 9]. The CG susceptibility $\chi_{\mathrm{CG}}$ is similarly expressed as
$\chi_{\mathrm{CG}}=\frac{3N}{C_{n}}\Bigg{[}\sum_{A>B}(q_{\kappa}^{AB})^{2}\Bigg{]}_{J},$ (7)
where
$q_{\kappa}^{AB}=\frac{1}{3N}\sum_{\alpha}^{3N}\kappa_{\alpha}^{(A)}\kappa_{\alpha}^{(B)},$ (8)
$\kappa_{\alpha}=\frac{1}{2\sqrt{2}}\sum_{\langle jk\rangle\in\alpha}J_{jk}\sin(\theta_{j}^{(A)}-\theta_{k}^{(A)}).$ (9)
Here, at each square plaquette $\alpha$ composed of nearest-neighbor bonds $\langle jk\rangle$, a local chirality variable $\kappa_{\alpha}$ is defined.
To estimate the critical temperature and some critical exponents of the 3D ±$\pm J$ XY model, we used the following dynamical scaling law for both SG and CG susceptibility[4, 9, 2]:
$\chi(t,T)=t^{\lambda(T)}\Phi\bigg{(}\frac{t}{\tau(T)}\bigg{)},$ (10)
where
$\lambda(T)=\frac{\gamma}{z_{\mathrm{s}}(T)\nu},$ (11)
$\tau(T)=(T-T_{\rm SG})^{-z_{\mathrm{s}}(T)\nu}.$ (12)
$z_{\mathrm{s}}(T)=z\frac{T_{\mathrm{SG}}}{T}.$ (13)
The relation expressed by Eq. (13) holds only within the context of the NER process in SG systems. Just above the transition temperature, Eq. (13) becomes equivalent to the definition of $z$. This relation was first noted for the Gaussian bond Ising model and gauge glass model[10]. It was later noted for Ising, XY, and Heisenberg SG systems[9]. The crossover between this dynamics and the normal dynamics is observed at the multicritical point in Ising SG systems[2]. However, it is noted that the temperature dependence of $z(T)$ has been confirmed for the SG case, but the validity of this relation in the CG case has not been sufficiently established over the entire temperature range. An analysis based on the aforementioned method has been previously reported[9]. Considering the advances in numerical techniques and available computational power since then, we conducted renewed simulations in the present study.
II.2 Observables and Scaling Law for Estimating the Dynamical Critical Exponent
In the scaling analysis based on the method described in Sec. II.1, the critical temperature $T_{\mathrm{c}}$ and the products of the critical exponents $\gamma/z\nu$ and $z\nu$ can be obtained. Therefore, to estimate all of the critical exponents, it is necessary to calculate $z$. On the basis of Eq. (1), we will explain the method for determining the dynamical correlation length $\xi(t)$ required to calculate $z$.
The SG correlation function $f_{\mathrm{SG}}(r)$ is defined by
$f_{\mathrm{SG}}(r)=\Bigg{[}\frac{1}{N}\sum_{i}^{N}\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{i+r}\rangle^{2}\Bigg{]}_{J},$ (14)
where $r$ denotes the distance between two spins in the lattice. This equation can be rewritten using Eq. (5) as
$f_{\mathrm{SG}}(r)=\frac{1}{NC_{n}}\Bigg{[}\sum_{A>B}\sum_{i,\mu,\nu}(q_{\mu\nu}^{AB}(i))(q_{\mu\nu}^{AB}(i+r))\Bigg{]}_{J},$ (15)
which is practically expressed as
$f_{\mathrm{SG}}(r)=\frac{1}{N}\Bigg{[}\sum_{i}^{N}$$\Bigg{\{}\Bigg{(}\frac{1}{n}\sum_{A=1}^{n}\boldsymbol{S}_{i}^{(A)}\cdot\boldsymbol{S}_{i+r}^{(A)}\Bigg{)}^{2}$ (16)
$-\frac{1}{n^{2}}\sum_{A=1}^{n}\Bigg{(}\boldsymbol{S}_{i}^{(A)}\cdot\boldsymbol{S}_{i+r}^{(A)}\Bigg{)}^{2}\Bigg{\}}\Bigg{]}_{J},$
for computations. This choice was made because of the computational complexity associated with Eq. (15), which requires computations in ${}_{n}C_{2}$ different ways.
Similarly, the CG correlation function $f_{\mathrm{CG}}$ is also defined by
$f_{\mathrm{CG}}$ $=\frac{1}{3N}\Bigg{[}\sum_{\alpha}^{3N}\Bigg{\{}\Bigg{(}\frac{1}{n}\sum_{A=1}^{n}\kappa_{\alpha}^{(A)}\kappa_{\alpha+r}^{(A)}\Bigg{)}^{2}$ (17)
$\qquad-\frac{1}{n^{2}}\sum_{A=1}^{n}\Bigg{(}\kappa_{\alpha}^{(A)}\kappa_{\alpha+r}^{(A)}\Bigg{)}^{2}\Bigg{\}}\Bigg{]}_{J}.$
For both SG and CG correlation functions, we assume the dynamical scaling ansatz[6, 7] expressed as
$\frac{f(r,t)}{\xi(t)^{-d+2-\eta_{\mathrm{eff}}}}=\Phi\bigg{(}\frac{r}{\xi(t)}\bigg{)}.$ (18)
Here, $d$ is the spatial dimension, $\eta_{\mathrm{eff}}$ is an effective scaling exponent, and $f(r,t)$ and $\xi(t)$ are the correlation function and correlation length, respectively, at Monte Carlo step (MCS) $t$. We estimate $\eta_{\mathrm{eff}}$ and dozens of $\xi(t)$ such that all of the $f(r,t)$ data collapse onto a single scaling function $\Phi(\cdot)$. Using the $f(r,t)$ computed at the estimated critical temperature and optimizing the parameters that satisfy the dynamical scaling ansatz Eq. (18), one can obtain sets of $\xi(t)$ at the critical temperature. In conjunction with Eq. (1), this approach enables the value of $z$ to be obtained. When evaluating $z$ in this context, the temperature $T$ in Eq. (13) is set to $T_{\mathrm{SG}}$.
When optimization calculations are performed, dozens of initial quantities $\xi(t)$ are required. For this purpose, we used $\xi(t)$ calculated by the second moment method. The SG and CG correlation length can be calculated by the second moment method as follows:
$\xi_{\mathrm{2nd,SG}}=\frac{1}{2\sin{(\pi/L)}}\sqrt{\frac{\tilde{\chi}_{\mathrm{SG}}(\mathbf{0})}{\tilde{\chi}_{\mathrm{SG}}(\mathbf{k}_{\mathrm{min}})}-1},$ (19)
$\xi_{\mathrm{2nd,CG}}=\frac{1}{2\sin{(\pi/L)}}\sqrt{\frac{\tilde{\chi}_{\mathrm{CG}}(\mathbf{0})}{\tilde{\chi}_{\mathrm{CG}}(\mathbf{k}_{\mathrm{min}})}-1},$ (20)
where $\tilde{\chi}_{\mathrm{SG}}(\cdot)$ and $\tilde{\chi}_{\mathrm{CG}}(\cdot)$ represent the Fourier transform of susceptibilities and $\mathbf{k}_{\mathrm{min}}$ denotes the smallest wave vector, which is determined as $\mathbf{k}_{\mathrm{min}}=(2\pi/L,0,0)$ in the present study.
The second moment correlation length can be calculated during the nonequilibrium relaxation process. However, the $z$ estimated using this method yields inaccurate values even for relatively simple models such as the 3D FM Ising model[5]. Therefore, an analysis based on Eq. (18) is necessary.
II.3 Gaussian Process Regression
In this subsection, we briefly explain the method of dynamic scaling analysis based on Gaussian process regression [11, 12, 6, 7]. This method is applied in all scaling analyses, and it is also used in Sec. IV.1 to assess the validity of the scaling plots. Let us write the scaling relation expressed in Eqs. (10) and (18) in the following form:
$Y=\Phi(X).$ (21)
We denote the data vectors formed by $X$ and $Y$ as $\vec{X}$ and $\vec{Y}$, respectively, and the dynamic scaling function evaluated at $\vec{X}$ as $\vec{\Phi}$. Then, the log-likelihood function for optimizable parameters $\vec{\theta}$ is given by
$\ln{\mathcal{L}(\vec{\theta})}=-\frac{1}{2}\ln{|2\pi\Sigma|}-\frac{1}{2}(\vec{Y}-\vec{\Phi})^{t}\Sigma^{-1}(\vec{Y}-\vec{\Phi}),$ (22)
where $\Sigma$ is the covariance matrix whose elements are defined via a kernel function $K(X_{i},X_{j},\vec{\theta})$ as
$\Sigma_{ij}=E_{i}^{2}\delta_{ij}+K(X_{i},X_{j},\vec{\theta}),$ (23)
$K(X_{i},X_{j},\vec{\theta})=\theta_{0}^{2}\delta_{ij}+\theta_{1}^{2}\exp\left(-\frac{(X_{i}-X_{j})^{2}}{2\theta_{2}^{2}}\right).$ (24)
Here, $E_{i}$ represents the error of $Y_{i}$, and $(\theta_{0},\theta_{1},\theta_{2})$ are the hyperparameters. By maximizing Eq. (22) using an optimization algorithm, the parameters can be estimated. Specifically, the dynamic scaling analysis using Eq. (10) yields the transition temperature and the exponents $\gamma/z\nu$ and $z\nu$, while the analysis using Eq. (18) provides the sets of $\xi(t)$ and $\eta_{\mathrm{eff}}$.
In the actual analysis, we applied a logarithmic transformation to the data to stabilize the fitting and improve numerical accuracy[12], defined as follows:
$X_{i}^{\prime}$ $\equiv\ln X_{i},$
$Y_{i}^{\prime}$ $\equiv\ln Y_{i},$
$E_{i}^{\prime}$ $\equiv\ln\left(1+\frac{E_{i}}{Y_{i}}\right)\approx\frac{E_{i}}{Y_{i}}.$ (25)
III Confirmation for the Estimation of the Dynamical Exponent
To calculate $z$, we first perform a scaling analysis based on Eq. (18) to obtain a set of dynamical correlation lengths $\xi(t)$ at the transition temperature and then estimate $z$ from the slope of $\xi(t)$ versus $t$ on a double-logarithmic plot according to Eq. (1). However, the reliability of obtained $z$ values using this method has not been thoroughly validated. Therefore, we first apply this method to calculate $z$ for the relatively well-studied 3D FM Ising model and the ±$\pm J$ Ising model to confirm if the results are consistent with previous studies and establish the reliability of this method. Following this confirmation, we apply the method to the primary focus of the present study, the 3D ±$\pm J$ XY model in the next section.
III.1 3D FM Ising Model
To analyze the 3D FM Ising model by the dynamical scaling method, we conducted simulations to 1000 MCS. Within this range, finite-size effects have been confirmed to be absent for $L\geq 159$[5]. We therefore adopted a linear size of $L=301$ in the present study. Under this condition, we computed the correlation function $f(r,t)$ at the critical temperature. Here, the sample number was set to 18,000 and the transition temperature was chosen as the highly accurate value obtained from a previous study, $T_{\mathrm{c}}=4.51152325$[13]. In addition, the MCSs for computing the correlation function were selected to be logarithmically spaced with 10 points between 10 MCS and 1000 MCS.
Figure 1 shows the dynamical correlation function of the 3D FM Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (18) to determine the set of correlation lengths required for an accurate estimation of $z$. In an optimization procedure, we discarded short-range correlations ($r<10$) and data of small $f(r,t)$ values ($f(r,t)<10^{-4}$). The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method. Figure 2 shows the scaling result based on the dynamical scaling law in Eq. (18). The resultant values of correlation length are plotted in Fig. 3 together with those obtained by the second moment method for comparison. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=2.018(8)$. In addition, the uncertainty was estimated as the standard deviation of the bootstrap distribution obtained from 10,000 resamplings[14]. This value of $z$ is consistent with the value $z=2.0235(8)$ obtained using the latest reliable methods[15], demonstrating the robustness and accuracy of our scaling analysis based on Eq. (18).
Figure 1: The dynamical correlation function of the 3D FM Ising model, $f(r,t)$, at the critical temperature $T_{\mathrm{C}}=4.51152325$[13] for various Monte Carlo steps $t$.
Figure 2: The scaling plot of the dynamical correlation function $f(r,t)$ for the 3D FM Ising model, based on the dynamical scaling law in Eq. (18). The quantities used in the scaling plot are the effective exponent $\eta_{\mathrm{eff}}=0.1433$ and the dozens of correlation length $\{\xi(t)\}$ shown in Fig. 3. These values were obtained through optimization calculations of Gaussian process regression.
Figure 3: The correlation length $\{\xi(t)\}$ obtained at the critical point of the 3D FM Ising model. Here, “inference” in the legend of the graph corresponds to $\{\xi(t)\}$ obtained through optimization calculations of the dynamical scaling law, whereas “second” represents $\{\xi(t)\}$ obtained using the second moment method. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
III.2 3D ±$\pm J$ Ising Model
Demonstrations of the calculation of $z$ for the 3D ±$\pm J$ Ising model were performed similarly as described in Sec. III.1. Simulations were conducted up to $5\times 10^{5}$ MCS. Within this range, finite-size effects were confirmed to vanish for $L\geq 31$[2]. In the present study, a linear size of $L=61$ was set, $n=256$ replicas were employed, and the sample number was set to 10. The critical temperature was chosen to be the highly accurate value obtained from a previous study, $T_{\mathrm{SG}}=1.178$[2]. The MC steps for computing the correlation function were selected to be logarithmically spaced with 50 points between 100 MCS and $5\times 10^{5}$ MCS.
Figure 4 shows the dynamical SG correlation function of the 3D ±$\pm J$ Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (18) to determine the set of correlation lengths required to accurately estimate $z$. During the optimization procedure, short-range correlations ($r<5$) and data with small $f(r,t)$ values ($f(r,t)<5\times 10^{-5}$) were excluded. The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method (Eq. (19)). Figure 5 shows the scaling result obtained using the dynamical scaling law in Eq. (18). The resultant values of correlation length are plotted in Fig. 6. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=5.92(3)$; its uncertainty, defined as the standard deviation of the bootstrap distribution, was assessed using 10 000 bootstrap trials[14]. This result is consistent with the value $z=5.89(3)$ obtained through extrapolation of the dynamical Binder parameter[2], demonstrating the robustness and accuracy of our scaling analysis approach based on Eq. (18).
Figure 4: The dynamical SG correlation function of the 3D ±$\pm J$ Ising model, $f_{\mathrm{SG}}(r,t)$, at the critical temperature $T_{\mathrm{SG}}=1.178$[2] for various Monte Carlo steps $t$.
Figure 5: The scaling plot of the dynamical SG correlation function $f_{\mathrm{SG}}(r,t)$ for the 3D ±$\pm J$ Ising model, based on the dynamical scaling law in Eq. (18). The quantities used in the scaling plot are the effective exponent $\eta_{\mathrm{eff}}=0.1773$ and the dozens of correlation length $\{\xi(t)\}$ shown in Fig. 6. These values were obtained through optimization calculations of Gaussian process regression.
Figure 6: The correlation length $\{\xi(t)\}$, as obtained through optimization calculations of the dynamical scaling law, at the critical point of the ±$\pm J$ Ising model in three dimensions. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
IV Results for 3D ±$\pm J$ XY model
To investigate the critical phenomena of SG systems using the dynamical scaling method, it is necessary to perform a scaling analysis based on Eq. (10); this analysis provides the transition temperature and the product of the critical exponents. For this purpose, identifying a linear size $L$ for which finite-size effects can be neglected within the range of considered MCSs is critical.
First, to examine size dependence, we investigated the behavior near the transition temperature of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ defined respectively by Eqs. (6) and (7). These physical quantities exhibit algebraic divergence at the transition temperature. According to Fig. 7, within the range of $5\times 10^{5}$ MCS, the size dependence of $\chi_{\mathrm{SG}}$ disappears for $L\geq 41$ and so does that of $\chi_{\mathrm{CG}}$ for $L\geq 51$. On the basis of these results, the size of the 3D ±$\pm J$ XY model was set to $L=91$ in the present study.
Figure 7: Size dependence of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ in the 3D ±$\pm J$ XY model: (a) $\chi_{\mathrm{SG}}$, where size dependence vanishes for $L\geq 41$, and (b) $\chi_{\mathrm{CG}}$, where size dependence vanishes for $L\geq 51$. The temperatures used for the calculations were $T=0.396$ for $\chi_{\mathrm{SG}}$ and $T=0.4460$ for $\chi_{\mathrm{CG}}$, which are close to the critical point.
Second, to estimate the transition temperature, we have to conduct a scaling analysis using the dynamical scaling law represented by Eq. (10). To perform this analysis, it is necessary to determine whether Eq. (13) can be applied to the dynamical scaling of the CG susceptibility. Therefore, we first examine the temperature dependence of $z$ at the CG transition. From the demonstrations described in Sec. III.1 and III.2, we note that the reliability of $z$ obtained from the correlation length estimated by the dynamical scaling law has been established. Based on this result, we calculated the CG correlation function over a temperature range that includes both low- and high-temperature regimes, using $n=256$ and 5 samples. Subsequently, a scaling analysis was conducted according to Eq. (18). In the optimization procedure, short-range correlations ($r<4$ for $f_{\mathrm{CG}}$) and data points with small values ($f_{\mathrm{CG}}(r,t)<2\times 10^{-6}$) were excluded. The initial values of $\{\xi(t)\}$ were set using the correlation length obtained via the second-moment method, as expressed in Eq. (20). The dynamical exponent $z$ was then estimated from the slope of the scaling function. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 10,000 resamplings[14]. The result is shown in Fig. 8. This figure clearly shows that, as in the SG case, the CG case also exhibits a dependence of $z$ on $1/T$. Accordingly, Eq. (13) should be employed in the dynamical scaling analysis of the CG susceptibility.
Figure 8: Temperature dependence of $z$ for the CG transition. The dashed line indicates the estimated CG transition temperature obtained in this study.
IV.1 Dynamical Scaling Analysis
Based on the previously obtained temperature dependence of the dynamical exponent $z(T)$, we performed a dynamical scaling analysis using the relation given in Eq. (10).
The quantities used in the calculations for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ were $n=256$, and 5 samples. The temperature range used for the calculations is shown in Fig. 9. In the scaling analysis for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$, the dynamical exponent $z$ was treated as temperature dependent, following the assumption described in Eq. (13). As a result, we obtained $T_{\mathrm{SG}}=0.396(1)$ and $T_{\mathrm{CG}}=0.4460(6)$. In addition, the product of critical exponents $\gamma/z\nu$ and $z\nu$ were estimated as scaling parameters. For $\chi_{\mathrm{SG}}$, $\gamma/z\nu=0.3660(4)$ and $z\nu=7.67(5)$ were obtained; for $\chi_{\mathrm{CG}}$, $\gamma/z\nu=0.3214(6)$ and $z\nu=5.70(1)$ were obtained. These results are compared with those reported in previous studies in Table 1 and Table 2.
The obtained transition temperature appeared to show good accuracy compared with previously reported values. The results confirmed that the SG transition temperature is lower than the CG transition temperature, supporting the spin chirality decoupling picture in SG systems [1].
Although the authors of some previous studies in which the NER method was used did not provide evidence supporting this theory [4, 9], our analysis of the transition temperature using a similar approach yielded results consistent with it. This discrepancy is attributable to finite-time corrections. The computational MCSs in the present study were approximately $5\times 10^{5}$ MCS, which is longer than the MCSs in previous studies[4, 9]. Figure 11 presents the relaxation data obtained in our study, scaled using the parameters from a previous study [9]. The inset shows the distribution of the log-likelihood values obtained via bootstrap sampling, comparing the fits using the parameters from a previous study [9] and those obtained in the present study. To evaluate the log-likelihood functions, we fixed the physical parameters such as $T_{\mathrm{SG}}$, $\gamma/z\nu$, and $z\nu$, and optimized the hyperparameters in the Gaussian process regression for each bootstrap sample. This procedure allows us to evaluate which set of physical parameters is more consistent with the dynamical scaling behavior of the relaxation data. In the prior study, physical quantities were estimated on the basis of relaxation data up to $1\times 10^{5}$ MCS. As shown in Fig. 11, the scaling was successful for data up to $1\times 10^{5}$ MCS. However, when the data were extended to $5\times 10^{5}$ MCS, the scaling failed, which is also clearly reflected in the behavior of the log-likelihood function. This discrepancy is likely attributable to finite-time corrections and supports the validity of our results. Similar issues have been pointed out in finite-size scaling analysis, where the picture can change for small systems[16, 17].
Figure 9: The results of the dynamical scaling analysis for the 3D ±$\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{SG}}$ and (b) the scaling plot. The scaling plot in (b) was obtained using the parameters $T_{\mathrm{SG}}=0.396(1)$, $\gamma/z\nu=0.3660(4)$, and $z\nu=7.67(5)$.
Figure 10: The results of the dynamical scaling analysis for the 3D ±$\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{CG}}$ and (b) the scaling plot. The scaling plot in (b) was generated using the parameters $T_{\mathrm{CG}}=0.4460(6)$, $\gamma/z\nu=0.3214(6)$, and $z\nu=5.70(1)$. In this scaling analysis, $z$ is treated as described in Eq. (13).
Figure 11: Scaling plots of the relaxation data obtained in our study, using the parameters from a previous study [9]: $T_{\mathrm{SG}}=0.435$, $\gamma/z\nu=0.368$, and $z\nu=6.25$. The insets show the distribution of the log-likelihood values obtained via bootstrap sampling, comparing the fits using the physical parameters from a previous study [9] and those obtained in the present study. The scaling plot in (a), using data up to $1\times 10^{5}$ MCS, shows a good data collapse. The scaling plot in (b), using data up to $5\times 10^{5}$ MCS, reveals that part of the data in the extended MCS range does not collapse onto the scaling function. The clearly separated distributions of the log-likelihood values show that the two scaling behaviors are distinguishable, which supports the physical conclusion drawn in this study.
IV.2 Estimation of critical exponents
Using the transition temperatures obtained from the analysis in Sec.IV.1, we evaluate the dynamical exponent $z$ at the critical temperatures. Based on these results and the products of critical exponents obtained in Sec.IV.1, we assess the critical behavior.
The results of the correlation function computed at the transition temperatures are shown in Fig. 12. We subsequently perform a dynamical scaling analysis based on Eq. (18). In an optimization procedure, we discarded short-range correlations ($r<3$ for $f_{\mathrm{SG}}$ and $r<4$ for $f_{\mathrm{CG}}$) and data of small values ($f_{\mathrm{SG}}(r,t)<10^{-5}$ and $f_{\mathrm{CG}}(r,t)<2\times 10^{-6}$). The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method, represented by Eq. (19) for SG and Eq. (20) for CG. The scaling plots by inference correlation length are shown in Fig. 13. Figure 14 shows the scaling result based on the dynamical scaling law in Eq. (18).
From the inference results and Eq. (1), the value of $z$ is evaluated as $z=5.86(1)$ for the SG case and $z=5.24(7)$ for the CG case. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 10,000 resamplings[14]. These results and the previously mentioned dynamical scaling analysis in Sec. IV.1 can be used to estimate the critical exponents. These results are compared with those of previous studies in Tables 1 and 2.
The obtained results show improved accuracy for both the transition temperature and the critical exponents compared with previously reported findings. Here, the value of $\gamma/z\nu$ is found to be almost the same as those reported in previous studies employing NER analysis for the SG transition, whereas this is not the case for the CG transition[4, 9]. This discrepancy is likely due to the fact that those earlier studies did not take into account the temperature dependence of the dynamical exponent $z$ in the CG analysis. The dynamical exponent $z$ was found to be lower in the CG case than in the SG case, which supports the result that finite-size effects are more pronounced in the CG case (Fig. 7).
Figure 12: The dynamical SG and CG correlation functions of the 3D ±$\pm J$ XY model at the critical temperature for various Monte Carlo steps $t$: (a) analysis of the SG transition and (b) analysis of the CG transition.
Figure 13: The scaling plots of the SG and CG correlation functions in the 3D ±$\pm J$ XY model: (a) analysis of the SG transition with $\eta_{\mathrm{eff}}=0.6865$, and (b) analysis of the CG transition with $\eta_{\mathrm{eff}}=1.169$.
Figure 14: The SG and CG correlation lengths in the 3D ±$\pm J$ XY model. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
Table 1: Analysis results of the SG transition in the 3D ±$\pm J$ XY model compared with the corresponding results reported in previous studies.
Reference $T_{\mathrm{SG}}$ $\gamma$ $z$ $\nu$
[18] 0.39(2) 4.4(3) 1.2(2)
[4] 0.455(15) 1.70(14) 5.6(6) 0.85(8)
[9] 0.435(15) 2.30(25) 5.0(1) 1.25(15)
Present work 0.396(1) 2.80(3) 5.86(1) 1.30(4)
Table 2: Analysis results of the CG transition in the 3D ±$\pm J$ XY model compared with those in previous studies.
Reference $T_{\mathrm{CG}}$ $\gamma$ $z$ $\nu$
[8] 0.39(3) 7.4(10) 1.2(2)
[4] 0.467(10) 1.41(8) 6.3(5) 0.74(4)
[9] 0.445(15) 1.65(2) 5.1(2) 1.05(10)
Present work 0.4460(6) 2.54(1) 5.24(7) 1.08(4)
V Summary and Discussion
We primarily aimed to validate a method for calculating the dynamical exponent from the correlation function, thereby enabling an analysis of SG systems. The obtained dynamical exponent appeared to be reliable, showed good accuracy, and was not limited to SG systems. The results confirm the reliability of this method. Using this approach, we analyzed the 3D ±$\pm J$ XY model. Compared with conventional methods, our method achieved a higher precision in evaluating the dynamical exponent. Through this analysis, we also examined the temperature dependence of $z(T)$ in the CG transition. In addition, our results support the validity of the spin chirality decoupling picture in SG systems.
Although some previous studies using the NER method did not support the spin chirality decoupling theory[4, 9], our analysis of the transition temperature using a similar method yielded results that support it. This discrepancy is attributable to the finite-time corrections. The computational MCS in this study was approximately $5\times 10^{5}$ MCS, which is longer than the MCSs in previous studies[4, 9], indicating the necessity of considering corrections in the analysis using the dynamical scaling method, which is a future research direction.
The correlations estimated using our method also face problems related to finite-time corrections. Because we do not extrapolate the scaling regime as $t\rightarrow\infty$ for the calculation of the dynamical exponent, our proposed method would suffer a difficult to apply to cases with strong fluctuations. An example of such a situation is the analysis of SG transitions near multi-critical points in the 3D ±$\pm J$ Ising model[2]. Making the method more widely applicable will require the development of an extrapolation method for $t\rightarrow\infty$. A reliable and systematic extrapolation scheme has recently been proposed in the NER method[19]. We expect that using such a scheme will lead to solutions to such problems. This work is also a future task.
Acknowledgments
This work was supported by JST SPRING, Grant Number JPMJSP2131. The authors are also grateful to the Supercomputer Center at the Institute for Solid State Physics, University of Tokyo, for use of their facilities.